# Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim2

### v3

@@ Abstract
We introduce SensorLLM, a two-stage framework that enables Large Language Models (LLMs) to perform human activity recognition (HAR) from wearable sensor data. While LLMs excel at reasoning and generalization, they struggle with time-series inputs due to limited semantic context, numerical complexity, and sequence variability. To address these challenges, we construct SensorQA, a question-answering dataset of human-intuitive sensor-text pairs spanning diverse HAR scenarios. It supervises the Sensor-Language Alignment stage, where the model aligns sensor inputs with trend descriptions. Special tokens are introduced to mark channel boundaries. This alignment enables LLMs to interpret numerical patterns, channel-specific signals, and variable-length inputs—without requiring human annotation. In the subsequent Task-Aware Tuning stage, we adapt the model for multivariate HAR classification, achieving performance that matches or exceeds state-of-the-art methods. Our results show that, guided by human-intuitive alignment, SensorLLM becomes an effective sensor learner, reasoner, and classifier—generalizing across varied HAR settings and paving the way for foundation model research in time-series analysis. Our codes are available at https://github.com/zechenli03/SensorLLM.
[Uncaptioned image] SensorLLM: Human-Intuitive Alignment of Multivariate Sensor Data with LLMs for Activity Recognition
Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim2 1University of New South Wales, Sydney 2University of Tokyo {zechen.li, s.deldari, hao.xue1, flora.salim}@unsw.edu.au chen-linyao217@g.ecc.u-tokyo.ac.jp
@@ 1 Introduction
Human Activity Recognition (HAR) is a time-series (TS) classification task that maps sensor signals, such as accelerometer and gyroscope data, to human activities. Traditional models like LSTM (Guan and Plötz, 2017; Hammerla et al., 2016) and DeepConvLSTM (Ordóñez and Roggen, 2016) learn high-level features but are task-specific and struggle to generalize across different sensor configurations and activity sets. In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Liu et al., 2023a; Wu et al., 2023b; Yin et al., 2023), including text and images.
@@
Enabling LLMs to process wearable sensor data (Jin et al., 2023) requires either (1) pretraining or fine-tuning on TS data (Zhou et al., 2023a), which demands substantial computational resources and is hindered by limited and imbalanced labeled data, or (2) leveraging zero-shot and few-shot prompting by converting sensor data into text (Kim et al., 2024; Ji et al., 2024). The latter approach avoids retraining but introduces key challenges: (i) Numerical encoding issues—language model tokenizers, designed for text, struggle with numerical values, treating consecutive numbers as independent tokens (Nate Gruver and Wilson, 2023) and failing to preserve temporal dependencies (Spathis and Kawsar, 2024). (ii) Sequence length constraints—complex numerical sequence often exceeds LLMs’ maximum context length, leading to truncation, information loss, and increased computational costs. (iii) Multivariate complexity—LLMs process univariate inputs, making it difficult to encode multivariate sensor data in a way that retains inter-channel dependencies. (iv) Prompt engineering challenges—designing effective prompts that enable LLMs to interpret numerical sensor readings, detect trends and classify activities remains a challenge (Liu et al., 2023b).
To address these challenges, we introduce SensorLLM, a human-intuition inspired framework that aligns numerical sensor data with natural language. Unlike image-text pairs, sensor data consists of multivariate signals with complex and varied patterns, making annotation particularly difficult. To tackle this, we release SensorQA, a question–answer (QA) dataset of aligned sensor-text pairs spanning diverse HAR tasks. SensorLLM is trained on this dataset to learn intuitive mappings between sensor readings and descriptive language, enabling LLMs to interpret sensor data through natural interactions—without any modification to the LLM architecture (see Figure 1).
@@
Existing methods (Jin et al., 2024a; Sun et al., 2024a) have explored condensed text prototypes for alignment, but these approaches often lack interpretability and require extensive tuning to select suitable prototypes. In contrast, we propose an automatic text generation approach that aligns with human intuition by deriving descriptive trend-based text directly from TS data using statistical analyses and predefined templates. This method is precise, scalable, and interpretable, eliminating the need for manual annotations while preserving essential sensor characteristics. SensorLLM follows a two-stage framework:
@@
Our SensorQA pairs each sensor window with QA pairs to align sensor features with natural language in a structured format. A pretrained TS encoder extracts temporal features from the sensor data, which are then projected into a space interpretable by the LLM, mitigating text-specific tokenization issues. We also introduce special tokens for each sensor channel, enabling LLMs to effectively capture multivariate (multi-channel) dependencies.
@@
The aligned embeddings are used for HAR, leveraging the LLM’s reasoning capabilities while keeping its parameters frozen. Importantly, our framework naturally supports sensor inputs of varying sequence lengths and arbitrary numbers of channels (i.e., multivariate TS data), a flexibility that prior approaches have struggled to accommodate.
@@
  • We release SensorQA, a novel human-intuitive dataset of aligned sensor-text QA pairs covering diverse HAR scenarios. We evaluate SensorLLM through text similarity metrics, human judgments, and LLM-based assessments, confirming its ability to capture temporal patterns for robust multimodal understanding. SensorQA and SensorLLM support sensor inputs with varying sequence lengths and channel configurations, allowing broad and realistic HAR evaluation.
  • SensorLLM achieves competitive results across five HAR datasets, matching or surpassing state-of-the-art models. Experiments further validate that modality alignment and task-specific prompts significantly enhance the LLM’s ability to interpret and classify sensor data.
  • We show that SensorLLM maintains strong performance in the Task-Aware Tuning Stage, even when applied to datasets distinct from those used during alignment, highlighting its robustness and generalizability in HAR tasks.
@@ 2 Related Work
In this section, we discuss recent developments in leveraging LLMs for time-series data, specifically focusing on two categories: (1) LLMs for time series as text and (2) Multimodal Large Language Models (MLLMs) for sensor data. A broader overview of other related works, including deep learning approaches to HAR and additional LLM-based forecasting methods, is provided in Appendix A.1.
@@
While LLMs excel in processing natural language, applying them directly to time-series data poses unique challenges (Spathis and Kawsar, 2024). Certain methods address this by treating time-series signals as raw text, using the same tokenization as natural language. Notable examples include PromptCast (Xue and Salim, 2023), which transforms numeric inputs into textual prompts for zero-shot forecasting, and LLMTime (Gruver et al., 2024), which encodes time-series as numerical strings for GPT-like models. However, due to the lack of specialized tokenizers for numeric sequences, LLMs may fail to capture crucial temporal dependencies and repetitive patterns (Spathis and Kawsar, 2024). To mitigate these issues, several works employ time-series encoders before mapping the resulting embeddings to language model spaces (Liu et al., 2024a; Zhou et al., 2023c; Xia et al., 2024), thus aligning sensor embeddings with textual embeddings in a contrastive or supervised manner.
@@ 3 Methods
In this work, we propose SensorLLM, a two-stage framework that aligns wearable sensor data with descriptive text by a high-precision corpus of question-answering pairs created without any human annotations and tailored for wearable sensor reasoning. Our aim is to build a multimodal model capable of interpreting and reasoning over time-series (TS) signals. As shown in Figure 2, SensorLLM consists of three core components: (1) a pretrained LLM, (2) a pretrained TS embedder, and (3) a lightweight MLP alignment module.
In the Sensor–Language Alignment stage, a generative model aligns sensor readings with text, and in the Task–Aware Tuning stage, a lightweight classifier is added on top of the LLM to perform HAR. Crucially, only the alignment MLP and this classifier are trainable—both the backbone LLM and the TS embedder remain frozen—resulting in just 5.67% (535.9 M) of parameters being fine-tuned in the first stage and 0.12% (10.5 M) in the second, making training extremely efficient.
@@
3.1 SensorQA Dataset
Aligning TS data with natural language is challenging due to the lack of rich semantic annotations beyond class labels, making manual labeling costly and impractical (Deldari et al., 2024; Haresamudram et al., 2024). While prior works often rely on fixed text prototypes (Sun et al., 2024b; Jin et al., 2024a), we introduce SensorQA —a scalable and human-intuitive dataset designed to bridge sensor data and language through structured QA pairs.
SensorQA is built on the idea that TS signals naturally exhibit semantic patterns—such as trends and statistical behaviors—that can be described in natural language. Using predefined templates (Appendix A.2), descriptive QA pairs are automatically generated for each sensor window without any human annotations. Each pair includes information such as sensor type, time range, and observed trends. Templates are randomly combined to enhance diversity. For example:
@@ 3.1 SensorQA Dataset
    The time-series data represents readings taken from a <S> sensor between <$t_{s}$> and <$t_{e}$> seconds.
@@
    To sum up, the data exhibited a <T> trend for a cumulative period of <$t_{t}$> seconds.
@@ 3.2 Sensor-Language Alignment
As shown in Figure 2 (a), the Sensor–Language Alignment stage uses a generative model to create multimodal sentences by combining single-channel sensor readings with text from our SensorQA dataset. The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels and $T$ is the sequence length. Each channel’s data, denoted as $\mathbf{X}^{c}$, is processed independently to retain channel-specific characteristics. The data is segmented into non-overlapping segments $\mathbf{X}_{S}^{c}$, where $S$ is the total number of segments. Each segment $x_{s}$ is assigned a random length $l$ within a predefined range, encouraging the model to learn from both short-term fluctuations and long-term trends.
@@
3.3 Task-Aware Tuning
Input Embedding.
For each sensor channel $c$, we retrieve its aligned embeddings $\hat{o}_{s}^{c}$ from the pretrained alignment module. These are concatenated across all channels, along with statistical features (mean and variance) from our SensorQA dataset, to form the final input embedding:
where $\hat{z}_{\text{stat}}$ represents the statistical information. This ensures the model integrates both temporal and statistical characteristics for HAR.
Loss Function.
$\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i},$ (4)
where $y_{i}$ is the ground truth label.
Method USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
PatchTST 45.2±$\pm 1.48$ 45.6±$\pm 2.19$ 86.8±$\pm 0.84$ 86.0±$\pm 0.71$ 82.0±$\pm 0.71$ 81.2±$\pm 0.84$ 80.0±$\pm 1.58$ 79.4±$\pm 1.34$ 35.6±$\pm 0.89$ 66.2±$\pm 1.10$
Ns-Transformer 52.6±$\pm 2.30$ 51.8±$\pm 2.86$ 88.0±$\pm 0.71$ 87.4±$\pm 0.55$ 78.8±$\pm 0.84$ 78.8±$\pm 0.84$ 77.2±$\pm 1.48$ 75.8±$\pm 1.48$ 34.8±$\pm 1.10$ 65.4±$\pm 0.55$
Informer 51.2±$\pm 1.30$ 51.6±$\pm 1.52$ 86.6±$\pm 1.14$ 86.4±$\pm 0.89$ 78.0±$\pm 1.58$ 78.6±$\pm 1.34$ 74.0±$\pm 0.71$ 72.8±$\pm 0.84$ 35.6±$\pm 0.55$ 66.8±$\pm 0.84$
Transformer 49.6±$\pm 1.67$ 50.6±$\pm 0.55$ 85.4±$\pm 0.89$ 85.2±$\pm 1.10$ 77.0±$\pm 0.71$ 77.6±$\pm 0.89$ 75.2±$\pm 1.30$ 74.6±$\pm 1.34$ 32.8±$\pm 0.84$ 65.4±$\pm 0.89$
iTransformer 48.4±$\pm 1.82$ 49.6±$\pm 1.67$ 81.8±$\pm 0.84$ 81.8±$\pm 0.84$ 76.6±$\pm 0.55$ 75.8±$\pm 0.45$ 80.4±$\pm 1.14$ 80.0±$\pm 1.22$ 19.8±$\pm 0.84$ 62.4±$\pm 0.89$
TimesNet 52.2±$\pm 2.39$ 52.6±$\pm 2.07$ 87.4±$\pm 1.14$ 86.6±$\pm 1.14$ 76.2±$\pm 1.92$ 77.4±$\pm 1.14$ 78.4±$\pm 1.52$ 77.2±$\pm 1.48$ 34.8±$\pm 0.84$ 65.8±$\pm 1.79$
GPT4TS 54.2±$\pm 2.05$ 56.0±$\pm 1.58$ 88.2±$\pm 0.84$ 87.6±$\pm 0.55$ 80.4±$\pm 0.89$ 79.8±$\pm 0.45$ 76.4±$\pm 1.14$ 75.4±$\pm 1.14$ 32.8±$\pm 1.10$ 62.2±$\pm 1.92$
Chronos+MLP 44.2±$\pm 1.30$ 44.0±$\pm 0.71$ 82.2±$\pm 0.84$ 81.2±$\pm 0.84$ 79.8±$\pm 0.45$ 79.8±$\pm 0.45$ 83.0±$\pm 0.71$ 82.0±$\pm 0.71$ 38.0±$\pm 0.71$ 68.2±$\pm 0.84$
DeepConvLSTM 48.8±$\pm 2.39$ 50.6±$\pm 2.41$ 89.2±$\pm 0.84$ 89.2±$\pm 0.84$ 78.4±$\pm 1.52$ 78.2±$\pm 1.10$ 75.0±$\pm 1.87$ 76.0±$\pm 1.00$ 40.4±$\pm 0.89$ 69.4±$\pm 1.14$
DeepConvLSTMAtt 54.0±$\pm 2.12$ 54.4±$\pm 3.21$ 89.6±$\pm 1.14$ 89.4±$\pm 1.14$ 79.2±$\pm 1.30$ 79.6±$\pm 1.14$ 77.4±$\pm 2.19$ 76.8±$\pm 1.48$ 41.4±$\pm 0.55$ 70.4±$\pm 0.55$
Attend 60.2±$\pm 2.17$ 60.8±$\pm 1.92$ 93.2±$\pm 0.84$ 92.8±$\pm 0.45$ 84.6±$\pm 1.14$ 85.0±$\pm 0.71$ 83.4±$\pm 1.14$ 82.6±$\pm 1.14$ 43.6±$\pm 0.55$ $\pm 0.71$
SensorLLM 61.2±$\pm 3.56$ 62.6±$\pm 3.36$ 91.2±$\pm 1.48$ 90.8±$\pm 1.30$ 86.2±$\pm 1.48$ 87.2±$\pm 0.84$ 89.4±$\pm 3.85$ 89.0±$\pm 3.54$ 48.6±$\pm 1.14$ 72.0±$\pm 0.71$

### v4

# Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim1

@@ Abstract
We introduce SensorLLM, a two-stage framework that enables Large Language Models (LLMs) to perform human activity recognition (HAR) from sensor time-series data. Despite their strong reasoning and generalization capabilities, LLMs remain underutilized for motion sensor data due to the lack of semantic context in time-series, computational constraints, and challenges in processing numerical inputs. SensorLLM addresses these limitations through a Sensor-Language Alignment stage, where the model aligns sensor inputs with trend descriptions. Special tokens are introduced to mark channel boundaries. This alignment enables LLMs to capture numerical variations, channel-specific features, and data of varying durations, without requiring human annotations. In the subsequent Task-Aware Tuning stage, we refine the model for HAR classification, achieving performance that matches or surpasses state-of-the-art methods. Our results demonstrate that SensorLLM evolves into an effective sensor learner, reasoner, and classifier through human-intuitive Sensor-Language Alignment, generalizing across diverse HAR datasets. We believe this work establishes a foundation for future research on time-series and text alignment, paving the way for foundation models in sensor data analysis. Our codes are available at https://github.com/zechenli03/SensorLLM.
[Uncaptioned image] SensorLLM: Aligning Large Language Models with Motion Sensors for Human Activity Recognition
Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim1 1University of New South Wales, Sydney 2University of Tokyo {zechen.li, s.deldari, hao.xue1, flora.salim}@unsw.edu.au chen-linyao217@g.ecc.u-tokyo.ac.jp
@@ 1 Introduction
Human Activity Recognition (HAR) is a time-series classification task that maps sensor signals, such as accelerometer and gyroscope data, to human activities. Traditional models like LSTM (Guan and Plötz, 2017; Hammerla et al., 2016) and DeepConvLSTM (Ordóñez and Roggen, 2016) learn high-level features but are task-specific and struggle to generalize across different sensor configurations and activity sets. In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Liu et al., 2023a; Wu et al., 2024; Yin et al., 2024), including text and images.
@@
Enabling LLMs to process wearable sensor data (Jin et al., 2023) requires either (1) pretraining or fine-tuning on TS data (Zhou et al., 2023a), which demands substantial computational resources and is hindered by limited and imbalanced labeled data, or (2) leveraging zero-shot and few-shot prompting by converting sensor data into text (Kim et al., 2024; Ji et al., 2024). The latter approach avoids retraining but introduces key challenges: (i) Numerical encoding issues. Language model tokenizers, designed for text, struggle with numerical values, treating consecutive numbers as independent tokens (Nate Gruver and Wilson, 2023) and failing to preserve temporal dependencies (Spathis and Kawsar, 2023). (ii) Sequence length constraints. Complex numerical sequence often exceeds LLMs’ maximum context length, leading to truncation, information loss, and increased computational costs. (iii) Multivariate complexity. LLMs process univariate inputs, making it difficult to encode multivariate sensor data in a way that retains inter-channel dependencies. (iv) Prompt engineering challenges. Designing effective prompts that enable LLMs to interpret numerical sensor readings, detect trends and classify activities remains a challenge (Liu et al., 2023b).
Unlike image-text pairs, sensor data comprises multi-channel numerical signals with diverse characteristics, making direct interpretation and annotation particularly difficult. Existing methods (Jin et al., 2024a; Sun et al., 2024) have explored condensed text prototypes for alignment, but these approaches often lack interpretability and require extensive tuning to select suitable prototypes.
@@
To address these challenges, we propose SensorLLM, a human-intuition inspired framework that aligns wearable sensor data with natural language without modifying the LLM itself. We propose an automatic text generation approach that aligns with human intuition by deriving descriptive trend-based text directly from time-series data using statistical analyses and predefined templates (see Figure 1). This method is precise, scalable, and interpretable, eliminating the need for manual annotations while preserving essential sensor characteristics. SensorLLM follows a two-stage framework:
@@
We automatically generate question-answer pairs to align sensor time-series with text while preserving temporal features using a pretrained encoder. The resulting embeddings are mapped into a space interpretable by the LLM, mitigating issues associated with text-specific tokenization. Additionally, we introduce special tokens for sensor channels, enabling LLMs to effectively capture multi-channel dependencies.
@@
The aligned embeddings are used for HAR, leveraging the LLM’s reasoning capabilities while keeping its parameters frozen. This design extends LLMs beyond their original training, addressing concerns raised by Tan et al. (2024) regarding their applicability to time-series data. Importantly, our framework naturally supports sensor inputs of varying sequence lengths and arbitrary numbers of channels (i.e., multivariate time-series), a flexibility that prior approaches have struggled to accommodate.
@@
  • We propose a human-intuitive approach for aligning sensor data with descriptive text, eliminating the need for manual annotations. SensorLLM support time-series inputs with varying sequence lengths and channel configurations, allowing broad and realistic HAR evaluation. We evaluate SensorLLM through text similarity metrics, human judgments, and LLM-based assessments, confirming its ability to capture temporal patterns for robust multimodal understanding.
  • SensorLLM achieves competitive results across five HAR datasets, matching or surpassing state-of-the-art models. Experiments further validate that modality alignment and task-specific prompts significantly enhance LLM’s ability to interpret and classify sensor data.
  • We show that SensorLLM maintains strong performance in the Task-Aware Tuning Stage, even when applied to datasets distinct from those used during alignment, highlighting its robustness and generalizability.
@@ 2 Related Work
In this section, we review recent advances in applying LLMs to time-series data, with a focus on three directions: (1) treating general time series as text for LLM-based modeling, (2) leveraging Multimodal Large Language Models (MLLMs) for sensor data, and (3) employing LLMs for Human Activity Recognition (HAR). A broader survey of related work, including deep learning approaches to HAR and additional LLM-based forecasting methods, is provided in Appendix A.1.
@@
While LLMs excel in processing natural language, applying them directly to time-series data poses unique challenges (Spathis and Kawsar, 2023). Certain methods address this by treating time-series signals as raw text, using the same tokenization as natural language. Notable examples include PromptCast (Xue and Salim, 2024), which transforms numeric inputs into textual prompts for zero-shot forecasting, and LLMTime (Gruver et al., 2023), which encodes time-series as numerical strings for GPT-like models. However, due to the lack of specialized tokenizers for numeric sequences, LLMs may fail to capture crucial temporal dependencies and repetitive patterns (Spathis and Kawsar, 2023). To mitigate these issues, several works employ time-series encoders before mapping the resulting embeddings to language model spaces (Liu et al., 2024a; Zhou et al., 2023b; Xia et al., 2024), thus aligning sensor embeddings with textual embeddings in a contrastive or supervised manner.
@@
LLMs for Human Activity Recognition.
Despite the success of LLMs in NLP, applying them to HAR remains challenging, as inertial signals are difficult to ground in language. Xia et al. (2023) introduce an unsupervised pipeline that employs a two-stage prompting scheme to infer activities from object sequences without manual descriptions. IMUGPT Leng et al. (2023) follows a data-synthesis approach: it prompts LLMs to generate diverse textual activity descriptions, converts them into 3D motion sequences, and then renders virtual IMU streams for training. HARGPT Ji et al. (2024) directly prompts LLMs to classify sensor time-series data, downsampling signals to 10Hz to enable zero-shot HAR from raw IMU inputs, and demonstrates that on simple tasks LLMs can match or even surpass traditional and deep learning models without domain-specific adaptation. ZARA Li et al. (2025) instead frames HAR as an agentic LLM workflow, supporting plug-and-play zero-shot motion time-series analysis with knowledge-guided reasoning and retrieval-enhanced evidence.
@@ 3 Methods
In this work, we propose SensorLLM, a two-stage framework that aligns wearable sensor data with descriptive text by high-precision question-answering pairs created without any human annotations and tailored for wearable sensor reasoning. Our aim is to build a multimodal model capable of interpreting and reasoning over time-series (TS) signals. As shown in Figure 2, SensorLLM consists of three core components: (1) a pretrained LLM, (2) a pretrained TS embedder, and (3) a lightweight MLP alignment module.
In the Sensor-Language Alignment Stage, a generative model aligns sensor readings with text , and in the Task–Aware Tuning stage, a lightweight classifier is added on top of the LLM to perform HAR. Crucially, only the alignment MLP and this classifier are trainable, while the backbone LLM and TS embedder remain frozen. This design results in just 5.67% (535.9 M) of parameters being fine-tuned in the first stage and 0.12% (10.5 M) in the second, making training extremely efficient.
@@
3.1 Sensor-Text Data Generation
Aligning time-series data with text for LLM-based tasks is challenging due to the lack of rich semantic labels beyond class annotations, making manual annotation impractical (Deldari et al., 2024; Haresamudram et al., 2025). While prior works rely on predefined text prototypes (Sun et al., 2024; Jin et al., 2024a), we aim for a more human-intuitive representation of sensor data.
We argue that time-series data inherently contains semantic patterns that can be expressed through descriptive text, from simple numerical trends to statistical insights. To achieve this, we automatically generate descriptive text by analyzing observed trends and fluctuations in the data. Using predefined templates, we construct diverse question-answer (QA) pairs that capture trend changes while ensuring accuracy and scalability. These templates (see Appendix A.2) are randomly combined to enhance diversity. For example:
@@ 3.1 Sensor-Text Data Generation
    The time-series data represents readings taken from a <S> sensor between $t_{s}$> and $t_{e}$> seconds.
@@
    To sum up, the data exhibited a <T> trend for a cumulative period of $t_{t}$> seconds.
@@ 3.2 Sensor-Language Alignment
As shown in Figure 2 (a), the Sensor-Language Alignment stage employs a generative model to create multimodal sentences that combine single-channel sensor readings with textual descriptions. The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels and $T$ is the sequence length. Each channel’s data, denoted as $\mathbf{X}^{c}$, is processed independently to retain channel-specific characteristics. The data is segmented into non-overlapping segments $\mathbf{X}_{S}^{c}$, where $S$ is the total number of segments. Each segment $x_{s}$ is assigned a random length $l$ within a predefined range, encouraging the model to learn from both short-term fluctuations and long-term trends.
@@
3.3 Task-Aware Tuning
Input Embedding.
For each sensor channel $c$, we retrieve its aligned sensor embeddings $\hat{o}_{s}^{c}$ from the pretrained alignment module. These embeddings are then concatenated across all channels, along with their corresponding statistical features (mean and variance), to form the final input embedding:
where $\hat{z}_{\text{stat}}$ represents the statistical information, and $C$ is the number of channels. This ensures the model integrates both temporal and statistical characteristics for HAR.
Loss Function.
$\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i},$ (4)
where $y_{i}$ is the ground truth label.